Present value.  A​ smooth-talking used-car salesman who smiles considerably is offering you a great deal on a​ "pre-owned" car. He​ says, "For only 7 annual payments of ​$2 comma 800​, this beautiful 1998 Honda Civic can be​ yours." If you can borrow money at 8​%, what is the price of this​ car? Assume the payment is made at the end of each year.
Answer #1

no. of annual payments ,n = 7

Annual payment = A = $2800

interest rate , r = 8% = 0.08

present value interest rate factor of annuity (PVIFA) =((1+r)n -1)/((1+r)n *r) = ((1.08)7 -1)/((1.08)7 *0.08) = 5.20637

price of car = A*PVIFA = 2800*5.20637 = $14,577.84